OpenStudy (anonymous):

Transform the function in vertex form 1. f(x)= x^2 + 8x - 6

OpenStudy (precal):

you will need to complete the square

OpenStudy (ash2326):

A parabola with vertex (h, k) has the equation \[y=a(x-h)^2+k\] We need to transform y=x^2-8x-6 in the above form. Can you try now?

OpenStudy (anonymous):

what does h and k stand for?

OpenStudy (ash2326):

(h, k) are the coordinates of the vertex of a parabola

OpenStudy (anonymous):

ok

OpenStudy (ash2326):

Can you try now?

OpenStudy (anonymous):

i still dont get it

OpenStudy (ash2326):

we have \[f(x)=x^2+8x-6\] and we need to convert it to \[f(x)=a(x-h)^2+k\] We need to find a, x and h, Let's try to find a We notice that in the equation \(f(x)=x^2+8x-6\), coefficient of x is 1 so \[a=1\] Do you get this?

OpenStudy (ash2326):

* coefficient of x^2

OpenStudy (anonymous):

You have to "complete the square" of f(x)=x^2+8x-6 \[f(x)=(x+4)^2-22\]So your vertex is at the point (-4,-22).

OpenStudy (anonymous):

Let me know if you are unsure how to do this

OpenStudy (anonymous):

Basically it works like this:\[f(x)=ax^2+bx+c\]transforms to\[f(x)=a(x-h)^2+k\]The coefficent on x^2 (=a) stays the same in both equations. To get the "h" in the second equation we take h=b/2. Now all that's left is the "k" value. Our k is found by noting that h^2+k=c...so k=c-h^2.

OpenStudy (precal):